OpenStudy (anonymous):

Can someone please help me with this problem ? Indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (anonymous):

the line to which it is parellel is y=2/3x +2... Since it is parellel, so the slope is same that is 2/3. Since the line contains the points 1,-2 the equation is : y+2=2/3(x-1) Now simplify.